The difference of a number and 20 is 16.
Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3816)


x = the unknown number
x-20 = difference of a number and 20
x-20 = 16
x = 16+20
x = 36

Answer by ikleyn(52754)

.

It means that the number is 20 units greater than 16, i.e. is 16 + 20 = 36.

It is how to solve the problem in 4-th grade without using equations.


(Because its level of complexity is the same as 1+2 = 3).